A 80.0-kg speedskater starts a race with an acceleration of 8.9 m/s^2. What is the net external force on him?
1
Expert's answer
2020-11-25T10:45:56-0500

We can find the net external force on the speedskater from the Newton's Second Law of Motion:


"F_{net}=ma=80.0\ kg\cdot\ 8.9\ \dfrac{m}{s^2}=712\ N."

Answer:

"F_{net}=712\ N."